The measurements of two complimentary angles in the ratio 1:5, Find the measurement of each angle from smallest to largest

Answer by checkley77(9427)

ADD THE RATIOS:
1+5=6
NOW DIVIDE THE TOTAL DEGREES BY THIS TOTAL:
180/6=30
THIS INDICATES THAT EACH UNIT OF THE RATIO HAS A VALUE OF 30.
THUS 1*30=30 FOR THE SMALLER ANGLE.
5*30=150 FOR THE LARGER ANGLE.
PROOFS:
30+150=180
-----------
1/5=30/150
150=150